The moon is 3.9 × 10⁵ km from Earth's center and 1.5 × 10⁸ km from the Sun's center. The masses of earth and the sun are 6.0 × 10²⁴ kg and 2.0 × 10³⁰ kg, respectively. Find the ratio of the gravitational fields due to Earth and the Sun at the center of the Moon.

1
Expert's answer
2021-11-18T10:17:05-0500

Explanation & Calculations


  • The field due to sun is

"\qquad\qquad
\begin{aligned}
\small E_s&=\small G\frac{m_s.m_m}{(1.5\times10^8\times10^3m)^2}\cdots (1)

\end{aligned}"

  • That due to earth is

"\qquad\qquad
\begin{aligned}
\small E_e&=\small G\frac{m_e.m_m}{(3.9\times10^5\times10^3m)^2}\cdots(2)
\end{aligned}"

  • By (2) "\small \div" (1), (ratio should be carefully determined from the question)

"\qquad\qquad
\begin{aligned}
\small E_e:E_s&=\small \frac{E_e}{E_s}\\
&=\small \frac{m_e}{m_s}\times\frac{(1.5\times10^{11})^2}{(3.9\times10^8)^2}\\
&=\small \frac{(6\times10^{24}kg)}{(2\times10^{30}kg)}\times\frac{(1.5\times10^{11})^2}
{(3.9\times10^8)^2}\\
&=\small 0.4
\end{aligned}"